OpenStudy (anonymous):

What is the sum of the squares to the solutions for the equation sqrt(1-cosx)+sqrt(1+cosx)=sqrt(3) that lies between -pi and pi?

OpenStudy (anonymous):

take \[\sqrt{1-cosx}=t\] we have \[\sqrt{1-t}=\sqrt{3}-\sqrt{1+t}\] now do squaring on both sides and find t

OpenStudy (anonymous):

you'll get x=pi/6 and 5pi/6

OpenStudy (anonymous):

So according to that, the answer is pi? That's very incorrect :(